sum of the series 222,224,226........888 is how much ? plz tell fast

Grade:10

2 Answers

Aryan Maurya
6 years ago
185370Its an APWhere a=222L=888D=224-222Calculate n by formula l= a +(n-1)dYou will get n=334 thenApplying formula of AP sumN/2(2a +(n-1)d) you get answer 185370
Vivek
6 years ago
Its an AP its have 334 terms. Now add by AP formula S=n/2(a+l). Its answer is 185370. You can find no of terms by simly subtraction and apply formula